OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x2 + 1,find f(g(3)).

OpenStudy (faiqraees):

direct answer or explanation?

OpenStudy (anonymous):

both please :)

OpenStudy (poohdessey):

you would plug in 3 for x in the g(x) equation

OpenStudy (anonymous):

i got 82?

OpenStudy (poohdessey):

therefore, g(x)= x2+1 g(3) = (3)2+1 g(3)= 6+1 g(3) = 7 i believe that is correct

OpenStudy (poohdessey):

is 7 an answer choice

rebeccaxhawaii (rebeccaxhawaii):

no that is incorrect

OpenStudy (anonymous):

no, it's 23,47,82, and 90.

OpenStudy (poohdessey):

hmm hold on

rebeccaxhawaii (rebeccaxhawaii):

first plug 3 into the g equation so g(x)=3 g(3)=3^2+1 g(3)=9+1 g(x)=10 then plug that into the f(x) equation so f(x)=10 f(10)=2(10)+3 F(10)=20+3 f(g(x))=23

rebeccaxhawaii (rebeccaxhawaii):

Please click "best response" for the person that helped you the most, this just shows the appreciation for the time they took to assist. If you want further help with anymore questions, please fan me so in the long run you can see when I'm on. I'll be sure to take a look at your question to help :)

OpenStudy (poohdessey):

right .. my bad

OpenStudy (anonymous):

yes! that's what I just got tysm! :)

rebeccaxhawaii (rebeccaxhawaii):

no problem